# Classical simulation of non-Gaussian bosonic circuits

### v1

@@ 3.1 Gaussian states
$\chi_{\rho}(\xi)=\operatorname{tr}\left(\rho D(\hat{d}^{-1}(\xi))\right)\qquad\text{for}\qquad\xi\in\mathbb{R}^{2n}\ ,$ (16)
@@
$D(\alpha)=e^{i\hat{d}(\alpha)^{T}\Omega R}\qquad\text{for}\qquad\alpha\in\mathbb{C}^{n}\ ,$ (17)
@@
[…]
@@
$d_{j}=\operatorname{tr}(\rho R_{j})\qquad\text{for}\qquad j\in[2n]\ .$ (21)
@@
$\operatorname{\Gamma}_{j,k}=\operatorname{tr}(\rho\{R_{j}-d_{j}I,R_{k}-d_{k}I\})\qquad\text{for}\qquad j,k\in[2n]$ (22)
are the second moments of the field operators. By definition, the covariance matrix is symmetric. It satisfies
@@
$\chi_{\rho}(\xi)=\exp\left(-\frac{1}{4}\xi^{T}\Omega^{T}\operatorname{\Gamma}\Omega\xi-d^{T}i\Omega\xi\right)\qquad\text{ with }\qquad\xi\in\mathbb{R}^{2n}\ .$ (24)
The pair $(\Gamma,d)$ are the covariance matrix and displacement of $\rho$. Since the characteristic function $\chi_{\rho}$ of a state $\rho$ fully determines the state $\rho$ by Eq. 20, a Gaussian state is fully determined by its first two moments. We will denote by $\rho(\operatorname{\Gamma},d)$ a Gaussian state with covariance matrix $\operatorname{\Gamma}$ and displacement vector $d$.
According to Williamson theorem [28] (see also [29, Appendix 6]), any symmetric matrix $\Gamma\in\mathsf{Mat}_{2n\times 2n}(\mathbb{R})$ can be written as
$\Gamma=S(D\oplus D)S^{T}\ ,$ (25)
where $D\in\mathsf{Mat}_{n\times n}(\mathbb{R})$ is a diagonal matrix and $S\in Sp(2n)$ where $Sp(2n)$ denotes the group of $2n\times 2n$ real symplectic matrices, i.e., $S$ satisfies $S\Omega S^{T}=\Omega$. Additionally, any symplectic matrix $S\in Sp(2n)$ can be decomposed as
@@
with $z_{j}>1,j\in[n]$. This is called the Euler or Bloch-Messiah decomposition [30]. Together with Eq. 25 this allows to express a covariance matrix $\Gamma$ of a pure state as
@@
displacement $D(\alpha)$ $e^{i\hat{d}(\alpha)^{T}\Omega R}$ $I$ $\hat{d}(\alpha)$
phase shift $F_{j}(\phi)$ $e^{-i\frac{\phi}{2}(Q_{j}^{2}+P_{j}^{2}-I)}$ $\begin{smallmatrix}&\begin{smallmatrix}Q_{j}\ \ \ &\ \ \ P_{j}\end{smallmatrix}\\ \begin{smallmatrix}Q_{j}\\ P_{j}\end{smallmatrix}&\left(\begin{smallmatrix}\cos(\phi)&\sin(\phi)\\ -\sin(\phi)&\cos(\phi)\\ \end{smallmatrix}\right)\end{smallmatrix}$ 00
beamsplitter $B_{j,k}(\omega)$ $e^{i\omega(Q_{j}Q_{k}+P_{j}P_{k})}$ $\begin{smallmatrix}&\begin{smallmatrix}\ \ \ Q_{j}\ \ \ &\ \ \ P_{j}\ \ \ &\ \ \ Q_{k}\ \ \ &\ \ \ P_{k}\ \ \ \end{smallmatrix}\\ \begin{smallmatrix}Q_{j}\\ P_{j}\\ Q_{k}\\ P_{k}\end{smallmatrix}&\left(\begin{smallmatrix}\cos(\omega)&0&0&\sin(\omega)\\ 0&\sin(\omega)&\cos(\omega)&0\\ 0&\cos(\omega)&-\sin(\omega)&0\\ -\sin(\omega)&0&0&\cos(\omega)\\ \end{smallmatrix}\right)\end{smallmatrix}$ 00
single-mode squeezing $S_{j}(z)$ $e^{i\frac{z}{2}(Q_{j}P_{j}+P_{j}Q_{j})}$ $\begin{smallmatrix}&\begin{smallmatrix}\ Q_{j}&P_{j}\end{smallmatrix}\ \\ \begin{smallmatrix}Q_{j}\\ P_{j}\end{smallmatrix}&\left(\begin{smallmatrix}e^{-z}&0\\ 0&e^{z}\\ \end{smallmatrix}\right)\end{smallmatrix}$ 00
Table 1: Unitary operator $U$, symplectic matrix $S\in Sp(2n)$ and vector $s\in\mathbb{C}^{n}$ associated with each Gaussian unitary operation. Apart from the case of the displacement, we only show the entries labelled by the associated field operators where $S$ is different from the identity.
@@ 3.2 Gaussian unitary operations
$UR_{j}U^{\dagger}=\sum_{j=1}^{2n}S_{j,k}R_{k}\qquad\text{with}\qquad S\in Sp(2n)\ .$ (33)
@@
$=\exp\left(-\hat{d}(\alpha)^{T}i\Omega\hat{d}(\beta)/2\right)D(\alpha+\beta)\qquad\text{for}\qquad\alpha,\beta\in\mathbb{C}^{n}\ .$ (38)
@@ 3.3 Coherent states
$D(\beta)|\alpha\rangle=|\alpha^{\prime}\rangle\qquad\text{with}\qquad\alpha^{\prime}=\exp(i\operatorname{Im}(\alpha^{T}\overline{\beta}))(\alpha-\beta)\in\mathbb{C}^{n}\qquad\text{for all}\qquad\alpha\in\mathbb{C}^{n}\ .$ (44)
For $\phi\in\mathbb{R}$ and $j\in[n]$, we have
$F_{j}(\phi)|\alpha\rangle=|\alpha^{\prime}\rangle\quad\text{with}\quad\alpha^{\prime}=(\alpha_{1},\ldots,\alpha_{j-1},e^{-i\phi}\alpha_{j},\alpha_{j+1},\ldots,\alpha_{n})\in\mathbb{C}^{n}\quad\text{for all}\quad\alpha\in\mathbb{C}^{n}$ (45)
and, for $\omega\in\mathbb{R}$ and $j,k\in[n]$,
@@
$\alpha^{\prime}_{k}$ $=-i\alpha_{j}\sin\omega+\alpha_{k}\cos\omega\ ,\qquad\qquad\text{for all}\qquad\alpha\in\mathbb{C}^{n}\ .$ (48)
$\alpha^{\prime}_{\ell}$ $=\alpha_{\ell}\qquad\text{for}\qquad\ell\notin\{j,k\}\ ,$ (49)
@@ 3.4 Gaussian measurements
We consider a $n$-mode Gaussian state $\rho\in\mathcal{B}(\mathcal{H}_{A}\otimes\mathcal{H}_{B})$ on a bipartite system where $\mathcal{H}_{A}\simeq\mathcal{H}_{k}$ includes $k$ modes and $\mathcal{H}_{B}\simeq H_{n-k}$ includes the remaining $n-k$ modes. The state $\rho_{AB}=\rho(V,s)$ has covariance matrix
$\operatorname{\Gamma}=\begin{pmatrix}\Gamma_{A}&\Gamma_{AB}\\ \Gamma_{AB}^{T}&\Gamma_{B}\end{pmatrix}\ ,$ (50)
@@
$p(\alpha)$ $=\frac{1}{\pi^{k}}\|\Pi_{\alpha}\rho_{AB}\|^{2}$ (53)
$=\frac{1}{\pi^{k}\sqrt{\det((\Gamma_{A}+I)/2)}}\exp\left(-(d(\alpha)-s_{A})^{T}(\Gamma_{A}+I)^{-1}(d(\alpha)-s_{A})\right)\ ,$ (54)
@@
The post-measurement state $\rho^{\prime}_{AB}=\rho(\operatorname{\Gamma}^{\prime},d^{\prime})$ is a Gaussian state with covariance matrix (see calculation e.g. in [31, Appendix B])
@@
$d^{\prime}=(\hat{d(\alpha)},s_{B}+\Gamma_{AB}^{T}(\Gamma_{A}+\Gamma_{M})^{-1}(\hat{d}(\alpha)-s_{A}))\ .$ (57)
@@ 3.5 Inner product formulas for Gaussian states
This allows us to compute the trace of two Gaussian states $\rho(\operatorname{\Gamma}_{1},d_{1})$ and $\rho(\operatorname{\Gamma}_{2},d_{2})$, with $\operatorname{\Gamma}_{1},\operatorname{\Gamma}_{2}\in\mathsf{Mat}_{2n\times 2n}(\mathbb{R})$ two covariance matrices and $d_{1},d_{2}\in\mathbb{R}^{2n}$. This amounts to evaluating a Gaussian integral, giving [27, 32]
$\operatorname{tr}\left(\rho(\operatorname{\Gamma}_{1},d_{1})\rho(\operatorname{\Gamma}_{2},d_{2})\right)=\frac{1}{\sqrt{\det\left((\operatorname{\Gamma}_{1}+\operatorname{\Gamma}_{2})/2\right)}}\exp\left(-(d_{2}-d_{1})^{T}(\operatorname{\Gamma}_{1}+\operatorname{\Gamma}_{2})^{-1}(d_{2}-d_{1})\right)\ .$ (59)
For pure states $\rho(\operatorname{\Gamma}_{1},d_{1})=|\psi_{1}\rangle\langle\psi_{1}|$ and $\rho(\operatorname{\Gamma}_{2},d_{2})=|\psi_{2}\rangle\langle\psi_{2}|$ this formula gives the inner product squared $|\langle\psi_{1},\psi_{2}\rangle|^{2}$. Notice that describing Gaussian states by their covariance matrix and displacement vector, e.g., using the characteristic function given in Eq. 24, does not allow computing the inner product $\langle\psi_{1},\psi_{2}\rangle$ (i.e., the inner product itself, not its absolute value). This is because the covariance matrix and displacement vector are functions of the density operator $\rho=|\psi\rangle\langle\psi|$ (recall the definitions in Eqs. 21 and 22), not of the state vector $|\psi\rangle$.
@@ 4.1 Extended description of a Gaussian state
[…]
@@
$\operatorname{Var}_{\psi}(R_{j})=\langle\psi,R_{j}^{2}\psi\rangle-\langle\psi,R_{j}\psi\rangle^{2}\qquad\text{for}\qquad j\in[2n]\ .$ (66)
@@
$\sup_{\alpha\in\mathbb{C}^{n}}|\langle\alpha,\psi\rangle|^{2}$ $=\max_{\alpha\in\mathbb{C}^{n}}|\langle\alpha,\psi\rangle|^{2}\geq\frac{1}{\sqrt{\frac{1}{2}\sum_{j=1}^{n}\left(\operatorname{Var}_{\psi}(Q_{j})+\operatorname{Var}_{\psi}(P_{j})-1\right)+1}}\ .$ (67)
@@
$\operatorname{Var}_{\psi}(Q_{j})\cdot\operatorname{Var}_{\psi}(P_{j})$ $\geq\frac{1}{4}\qquad\text{for}\qquad j\in[n]\ .$ (68)
@@
$\frac{1}{2}(a+b)\geq\sqrt{ab}\qquad\text{for}\qquad a,b\in\mathbb{R}$ (69)
@@
$\operatorname{Var}_{\psi}(Q_{j})+\operatorname{Var}_{\psi}(P_{j})\geq 1\qquad\text{for}\qquad j\in[n]\ .$ (70)
@@ 4.2 Computing the overlap between Gaussian states
In this section we give an algorithm $\mathsf{overlap}$ which computes overlaps between Gaussian states. This algorithm uses an expression given in Lemma 4.4 which relates inner products between three Gaussian states and a displacement operator $D(\alpha),\alpha\in\mathbb{C}^{n}$. This is a generalization of a formula derived in Ref. [27] for the case of centered states and $\alpha$ set to zero.
@@
$=\frac{\exp\left(-{d_{1}}^{T}W_{1}{d_{1}}-{d_{2}}^{T}W_{2}{d_{2}}+d_{1}^{T}W_{3}d_{2}-\hat{d}(\alpha)^{T}\Omega^{T}\operatorname{\Gamma}_{5}\Omega\hat{d}(\alpha)-i\hat{d}(\alpha)^{T}\Omega^{T}(W_{4}d_{1}+W_{5}d_{2}+d_{3})\right)}{\sqrt{\det(\Omega^{T}(\operatorname{\Gamma}_{2}+\operatorname{\Gamma}_{3})\Omega/2)\det(\Omega^{T}(\operatorname{\Gamma}_{1}+\operatorname{\Gamma}_{4})\Omega/2)}}$ (76)
@@
For convenience, we introduce an algorithmic routine $\mathsf{overlaptriple}$ which evaluates Eq. 76 in Lemma 4.4. This routine is one of the main ingredients in the algorithms $\mathsf{overlap}$ and $\mathsf{postmeasure}$. The algorithm $\mathsf{overlaptriple}$ takes covariance matrices $\operatorname{\Gamma}_{1},\operatorname{\Gamma}_{1},\operatorname{\Gamma}_{2}$ and displacement vector $d_{1},d_{2},d_{3}$ of three Gaussian states $\psi_{1},\psi_{2},\psi_{3}$ as well as overlaps $u=\langle\psi_{3},D(\lambda)\psi_{1}\rangle,v=\langle\psi_{1},\psi_{2}\rangle$ and computes the overlap $\langle\psi_{2},\psi_{3}\rangle$. It has runtime $O(n^{3})$ as it consists in basic linear algebra operations. We give pseudocode for the algorithm in Fig. 1.
@@
2:$o\leftarrow u^{-1}v^{-1}T$triangleright compute the overlap $\langle\psi_{2},\psi_{3}\rangle$
@@
Figure 1: The algorithm $\mathsf{overlaptriple}$ takes as input the covariance matrices $\operatorname{\Gamma}_{j}$ and displacement vectors $d_{j}$ of three pure Gaussian states $|\psi_{j}\rangle,j\in[3]$ and the overlaps $\langle\psi_{3},D(\lambda)\psi_{1}\rangle,\langle\psi_{1},\psi_{2}\rangle$. The latter have to be non-zero. The algorithm computes the overlap $\langle\psi_{2},\psi_{3}\rangle$ using Eq. 76 in Lemma 4.4.
(a) The input to the algorithm $\mathsf{overlaptriple}$ consists of covariance matrices and displacements of three Gaussian states $\psi_{1},\psi_{2},\psi_{3}$ together with non-zero overlaps $u=\langle\psi_{3},D(\lambda)\psi_{1}\rangle,v=\langle\psi_{1},\psi_{2}\rangle$.
(b) Applying $\mathsf{overlaptriple}$ provides the inner product $\langle\psi_{2},\psi_{3}\rangle$. In this diagrammatic representation, this completes the triangle with vertices $\psi_{1},\psi_{2},\psi_{3}$.
Figure 2: A graphical representation of the algorithm $\mathsf{overlaptriple}$. Solid lines represent inner products between the states at the vertices that are given / have been computed and are non-zero. Inner products of the form $\langle\psi_{3},D(\lambda)\psi_{1}\rangle$ are represented by arrows.
We give an efficient algorithm $\mathsf{overlap}$ which computes inner products $\langle\psi(\Delta_{1}),\psi(\Delta_{2})\rangle$ from descriptions $(\Delta_{1},\Delta_{2})\in\operatorname{\mathsf{Desc}}_{n}\times\operatorname{\mathsf{Desc}}_{n}$. We give pseudocode for the algorithm in Fig. 3, we show the associated claims in Theorem 4.5 and we illustrate the algorithm in Fig. 4.
@@
1:$(\operatorname{\Gamma}^{\prime}_{1},\operatorname{\Gamma}^{\prime}_{2},\operatorname{\Gamma}^{\prime}_{3})\leftarrow(I,\operatorname{\Gamma}_{1},\operatorname{\Gamma}_{2})$triangleright covariance matrices of $(|\alpha_{1}\rangle,\psi(\operatorname{\Delta}_{1}),\psi(\operatorname{\Delta}_{2}))$
$(d^{\prime}_{1},d^{\prime}_{2},d^{\prime}_{3})\leftarrow(\hat{d}(\alpha_{1}),\hat{d}(\alpha_{1}),\hat{d}(\alpha_{2}))$triangleright displacement vectors of $(|\alpha_{1}\rangle,\psi(\operatorname{\Delta}_{1}),\psi(\operatorname{\Delta}_{2}))$
3:$\lambda\leftarrow\alpha_{1}-\alpha_{2}$
4:$u\leftarrow e^{-i\operatorname{Im}(\alpha_{1}^{T}\overline{\alpha_{2}})}\overline{r_{2}}$triangleright $u=\langle\psi(\Delta_{2}),D(\lambda)\alpha_{1}\rangle$
5:$v\leftarrow r_{1}$triangleright $v=\langle\alpha_{1},\psi(\operatorname{\Delta}_{1})\rangle$
6:$w\leftarrow\mathsf{overlaptriple}(\operatorname{\Gamma}^{\prime}_{1},\operatorname{\Gamma}^{\prime}_{2},\operatorname{\Gamma}^{\prime}_{3},d^{\prime}_{1},d^{\prime}_{2},d^{\prime}_{3},u,v,\lambda)$triangleright compute  $\langle\psi(\operatorname{\Delta}_{1}),\psi(\operatorname{\Delta}_{2})\rangle$
@@
Figure 3: The algorithm $\mathsf{overlap}$ takes descriptions $\Delta_{1},\Delta_{2}\in\operatorname{\mathsf{Desc}}_{n}$ and outputs the overlap $\langle\psi(\Delta_{1}),\psi(\Delta_{2})\rangle$.
@@
The algorithm $\mathsf{overlap}:\operatorname{\mathsf{Desc}}_{n}\times\operatorname{\mathsf{Desc}}_{n}\rightarrow\mathbb{C}$ given in Fig. 3 runs in time $O(n^{3})$. It satisfies
$\mathsf{overlap}(\operatorname{\Delta}_{1},\operatorname{\Delta}_{2})$ $=\langle\psi(\operatorname{\Delta}_{1}),\psi(\operatorname{\Delta}_{2})\rangle\qquad\textrm{ for all }\qquad\operatorname{\Delta}_{1},\operatorname{\Delta}_{2}\in\operatorname{\mathsf{Desc}}_{n}\ .$ (86)

### v2

@@ 2.1 Gaussian states
$\chi_{\rho}(\xi)=\operatorname{tr}\left(\rho D(\hat{d}^{-1}(\xi))\right)\quad\text{for}\quad\xi\in\mathbb{R}^{2n}\ ,$ (19)
@@
$D(\alpha)=e^{i\hat{d}(\alpha)^{T}\Omega R}\quad\text{for}\quad\alpha\in\mathbb{C}^{n}\ ,$ (20)
@@
$\begin{gathered}\begin{aligned} \hat{d}:\mathbb{C}^{n}&\rightarrow\mathbb{R}^{2n}\\ \alpha&\mapsto\hat{d}(\alpha)=\sqrt{2}\left(\operatorname{Re}(\alpha_{1}),\operatorname{Im}(\alpha_{1}),\ldots,\operatorname{Re}(\alpha_{n}),\operatorname{Im}(\alpha_{n})\right)\end{aligned}\end{gathered}\ .$ (22)
@@
$d_{j}=\operatorname{tr}(\rho R_{j})\quad\text{for}\quad j\in[2n]\ .$ (24)
@@
$\operatorname{\Gamma}_{j,k}=\operatorname{tr}(\rho\{R_{j}-d_{j}I,R_{k}-d_{k}I\})$ (25)
for $j,k\in[2n]$ are the second moments of the field operators. By definition, the covariance matrix is symmetric. It satisfies
@@
$\chi_{\rho}(\xi)=\exp\left(-\frac{1}{4}\xi^{T}\Omega^{T}\operatorname{\Gamma}\Omega\xi-d^{T}i\Omega\xi\right)$ (27)
with $\xi\in\mathbb{R}^{2n}$. The pair $(\Gamma,d)$ are the covariance matrix and displacement of $\rho$. Since the characteristic function $\chi_{\rho}$ of a state $\rho$ fully determines the state $\rho$ by Eq. 23, a Gaussian state is fully determined by its first two moments. We will denote by $\rho(\operatorname{\Gamma},d)$ a Gaussian state with covariance matrix $\operatorname{\Gamma}$ and displacement vector $d$.
According to Williamson theorem [a6cda7d0-29df-361a-a64d-8dd6c762bce6] (see also [arnol2013mathematical, Appendix 6]), any symmetric matrix $\Gamma\in\mathsf{Mat}_{2n\times 2n}(\mathbb{R})$ can be written as (see e.g., [degossonSymplecticGeometryQuantum2006, Theorem 8.11])
$\Gamma=S\left(\bigoplus_{j=1}^{n}\begin{pmatrix}d_{j}&0\\ 0&d_{j}\end{pmatrix}\right)S^{T}\ ,$ (28)
where $d_{j}\in\mathbb{R}$ for $j\in[n]$ and $S\in Sp(2n)$ where $Sp(2n)$ denotes the group of $2n\times 2n$ real symplectic matrices, i.e., $S$ satisfies $S\Omega S^{T}=\Omega$. Additionally, any symplectic matrix $S\in Sp(2n)$ can be decomposed as
@@
with $z_{j}>1,j\in[n]$. This is called the Euler or Bloch-Messiah decomposition [Arvind_1995]. Together with Eq. 28, the Euler decomposition allows to express a covariance matrix $\Gamma$ of a pure state as
@@
displacement $D(\alpha)$ $e^{i\hat{d}(\alpha)^{T}\Omega R}$ II $\hat{d}(\alpha)$
phase shift $F_{j}(\phi)$ $e^{-i\frac{\phi}{2}(Q_{j}^{2}+P_{j}^{2}-I)}$ $\begin{matrix}&\begin{matrix}Q_{j}\ \ \ &\ \ \ P_{j}\end{matrix}\\ \begin{matrix}Q_{j}\\ P_{j}\end{matrix}&\left(\begin{matrix}\cos(\phi)&\sin(\phi)\\ -\sin(\phi)&\cos(\phi)\\ \end{matrix}\right)\end{matrix}$ 0
beamsplitter $B_{j,k}(\omega)$ $e^{i\omega(Q_{j}Q_{k}+P_{j}P_{k})}$ $\begin{matrix}&\begin{matrix}\ \ \ Q_{j}\ \ \ &\ \ \ P_{j}\ \ \ &\ \ \ Q_{k}\ \ \ &\ \ \ P_{k}\ \ \ \end{matrix}\\ \begin{matrix}Q_{j}\\ P_{j}\\ Q_{k}\\ P_{k}\end{matrix}&\left(\begin{matrix}\cos(\omega)&0&0&\sin(\omega)\\ 0&\sin(\omega)&\cos(\omega)&0\\ 0&\cos(\omega)&-\sin(\omega)&0\\ -\sin(\omega)&0&0&\cos(\omega)\\ \end{matrix}\right)\end{matrix}$ 0
single-mode squeezing $S_{j}(z)$ $e^{i\frac{z}{2}(Q_{j}P_{j}+P_{j}Q_{j})}$ $\begin{matrix}&\begin{matrix}\ Q_{j}&P_{j}\end{matrix}\ \\ \begin{matrix}Q_{j}\\ P_{j}\end{matrix}&\left(\begin{matrix}e^{-z}&0\\ 0&e^{z}\\ \end{matrix}\right)\end{matrix}$ 0
[…]
@@ 2.2 Gaussian unitary operations
$UR_{j}U^{\dagger}=\sum_{j=1}^{2n}S_{j,k}R_{k}\quad\text{with}\quad S\in Sp(2n)\ .$ (36)
@@
$=\exp\left(-\hat{d}(\alpha)^{T}i\Omega\hat{d}(\beta)/2\right)D(\alpha+\beta)$ (41)
for $\alpha,\beta\in\mathbb{C}^{n}$.
@@ 2.3 Coherent states
$D(\beta)|\alpha\rangle=|\alpha^{\prime}\rangle$ (47)
with $\alpha^{\prime}=\exp(i\operatorname{Im}(\alpha^{T}\overline{\beta}))(\alpha-\beta)\in\mathbb{C}^{n}$ for all $\alpha\in\mathbb{C}^{n}$. For $\phi\in\mathbb{R}$ and $j\in[n]$, we have
$F_{j}(\phi)|\alpha\rangle=|\alpha^{\prime}\rangle$ (48)
with $\alpha^{\prime}=(\alpha_{1},\ldots,\alpha_{j-1},e^{-i\phi}\alpha_{j},\alpha_{j+1},\ldots,\alpha_{n})\in\mathbb{C}^{n}$ for all $\alpha\in\mathbb{C}^{n}$ and, for $\omega\in\mathbb{R}$ and $j,k\in[n]$,
@@
$\alpha^{\prime}_{k}$ $=-i\alpha_{j}\sin\omega+\alpha_{k}\cos\omega\ ,\quad\text{for all}\quad\alpha\in\mathbb{C}^{n}\ .$ (51)
$\alpha^{\prime}_{\ell}$ $=\alpha_{\ell}\quad\text{for}\quad\ell\notin\{j,k\}\ ,$ (52)
@@ 2.4 Gaussian measurements
We consider a $n$-mode Gaussian state $\rho\in\mathcal{B}(\mathcal{H}_{A}\otimes\mathcal{H}_{B})$ on a bipartite system where $\mathcal{H}_{A}\simeq\mathcal{H}_{k}$ includes $k$ modes and $\mathcal{H}_{B}\simeq H_{n-k}$ includes the remaining nkn-k modes. The state $\rho_{AB}=\rho(V,s)$ has covariance matrix
$\operatorname{\Gamma}=\begin{pmatrix}\Gamma_{A}&\Gamma_{AB}\\ \Gamma_{AB}^{T}&\Gamma_{B}\end{pmatrix}\$, (53)
@@
$p(\alpha)$ $=\frac{1}{\pi^{k}}\operatorname{tr}(\Pi_{\alpha}\rho_{AB})$ (56)
$=\frac{\exp\left(-(d(\alpha)-s_{A})^{T}(\Gamma_{A}+I)^{-1}(d(\alpha)-s_{A})\right)}{\pi^{k}\sqrt{\det(\Gamma_{A}+I)}}\ ,$ (57)
@@
The post-measurement state $\rho^{\prime}_{AB}=\rho(\operatorname{\Gamma}^{\prime},d^{\prime})$ is a Gaussian state with covariance matrix (see calculation e.g. in [garcia2007quantum, Appendix B])
@@
$d^{\prime}=(\hat{d(\alpha)},s_{B}+\Gamma_{AB}^{T}(\Gamma_{A}+I)^{-1}(\hat{d}(\alpha)-s_{A}))\ .$ (60)
@@ 2.5 Inner product formulas for Gaussian states
This allows us to compute the trace of two Gaussian states $\rho(\operatorname{\Gamma}_{1},d_{1})$ and $\rho(\operatorname{\Gamma}_{2},d_{2})$, with $\operatorname{\Gamma}_{1},\operatorname{\Gamma}_{2}\in\mathsf{Mat}_{2n\times 2n}(\mathbb{R})$ two covariance matrices and $d_{1},d_{2}\in\mathbb{R}^{2n}$. This amounts to evaluating a Gaussian integral, giving [PhysRevA.61.022306, PhysRevLett.115.260501]
$\operatorname{tr}\left(\rho(\operatorname{\Gamma}_{1},d_{1})\rho(\operatorname{\Gamma}_{2},d_{2})\right)$ (62)
$=\frac{\exp\left(-(d_{2}-d_{1})^{T}(\operatorname{\Gamma}_{1}+\operatorname{\Gamma}_{2})^{-1}(d_{2}-d_{1})\right)}{\sqrt{\det\left((\operatorname{\Gamma}_{1}+\operatorname{\Gamma}_{2})/2\right)}}\ .$ (63)
For pure states $\rho(\operatorname{\Gamma}_{1},d_{1})=|\psi_{1}\rangle\langle\psi_{1}|$ and $\rho(\operatorname{\Gamma}_{2},d_{2})=|\psi_{2}\rangle\langle\psi_{2}|$ this formula gives the inner product squared $|\langle\psi_{1},\psi_{2}\rangle|^{2}$. Notice that describing Gaussian states by their covariance matrix and displacement vector, e.g., using the characteristic function given in Eq. 27, does not allow computing the inner product $\langle\psi_{1},\psi_{2}\rangle$ (i.e., the inner product itself, not its absolute value). This is because the covariance matrix and displacement vector are functions of the density operator $\rho=|\psi\rangle\langle\psi|$ (recall the definitions in Eqs. 24 and 25), not of the state vector $|\psi\rangle$.
@@ 3.1 Extended description of a Gaussian state
$\begin{gathered}\begin{aligned} \psi:\quad\mathsf{Desc}_{n}\quad&\rightarrow\quad\mathrm{Gauss}_{n}\\ \Delta\quad&\mapsto\quad\psi(\Delta)\end{aligned}\end{gathered}\ .$ (69)
@@
$\operatorname{Var}_{\psi}(R_{j})=\langle\psi,R_{j}^{2}\psi\rangle-\langle\psi,R_{j}\psi\rangle^{2}\quad\text{for}\quad j\in[2n]\ .$ (70)
@@
$\sup_{\alpha\in\mathbb{C}^{n}}|\langle\alpha,\psi\rangle|^{2}=\max_{\alpha\in\mathbb{C}^{n}}|\langle\alpha,\psi\rangle|^{2}$ (71)
$\qquad\geq\frac{1}{\sqrt{\frac{1}{2}\sum_{j=1}^{n}\left(\operatorname{Var}_{\psi}(Q_{j})+\operatorname{Var}_{\psi}(P_{j})-1\right)+1}}\ .$ (72)
@@
$\operatorname{Var}_{\psi}(Q_{j})\cdot\operatorname{Var}_{\psi}(P_{j})$ $\geq\frac{1}{4}\quad\text{for}\quad j\in[n]\ .$ (73)
@@
$\frac{1}{2}(a+b)\geq\sqrt{ab}\quad\text{for}\quad a,b\in\mathbb{R}$ (74)
@@
$\operatorname{Var}_{\psi}(Q_{j})+\operatorname{Var}_{\psi}(P_{j})\geq 1\quad\text{for}\quad j\in[n]\ .$ (75)
@@ 3.2 Computing the overlap between Gaussian states
In this section we give an algorithm $\mathsf{overlap}$ which computes overlaps between Gaussian states. This algorithm uses an expression given in Lemma 3.4 which relates inner products between three Gaussian states and a displacement operator $D(\alpha),\alpha\in\mathbb{C}^{n}$. This is a generalization of a formula derived in Ref. [PhysRevA.61.022306] for the case of centered states and $\alpha$ set to zero.
@@
$=\frac{\exp\left(-{d_{1}^{\prime}}^{T}W_{1}{d_{1}^{\prime}}-{d_{2}^{\prime}}^{T}W_{2}{d_{2}^{\prime}}+{d_{1}^{\prime}}^{T}W_{3}d_{2}^{\prime}-\hat{d}(\alpha)^{T}\Omega^{T}\operatorname{\Gamma}_{5}\Omega\hat{d}(\alpha)-i\hat{d}(\alpha)^{T}\Omega^{T}(W_{4}d_{1}^{\prime}+W_{5}d_{2}^{\prime}+d_{3})\right)}{\sqrt{\det(\Omega^{T}(\operatorname{\Gamma}_{2}+\operatorname{\Gamma}_{3})\Omega/2)\det(\Omega^{T}(\operatorname{\Gamma}_{1}+\operatorname{\Gamma}_{4})\Omega/2)}}$ (81)
@@
For convenience, we introduce an algorithmic routine $\mathsf{overlaptriple}$ which evaluates Eq. 81 in Lemma 3.4. This routine is one of the main ingredients in the algorithms $\mathsf{overlap}$ and $\mathsf{postmeasure}$. The algorithm $\mathsf{overlaptriple}$ takes covariance matrices $\operatorname{\Gamma}_{1},\operatorname{\Gamma}_{1},\operatorname{\Gamma}_{2}$ and displacement vector $d_{1},d_{2},d_{3}$ of three Gaussian states $\psi_{1},\psi_{2},\psi_{3}$ as well as overlaps $u=\langle\psi_{3},D(\lambda)\psi_{1}\rangle,v=\langle\psi_{1},\psi_{2}\rangle$ and computes the overlap $\langle\psi_{2},\psi_{3}\rangle$. It has runtime $O(n^{3})$ as it consists of basic linear algebra operations. We give pseudocode for the algorithm.
@@
[…]
@@
[…]
We give an efficient algorithm $\mathsf{overlap}$ which computes inner products $\langle\psi(\Delta_{1}),\psi(\Delta_{2})\rangle$ from descriptions $(\Delta_{1},\Delta_{2})\in\operatorname{\mathsf{Desc}}_{n}\times\operatorname{\mathsf{Desc}}_{n}$. We give pseudocode for the algorithm, illustrate it in Fig. 4 and show the associated claims in Theorem 3.5.
@@
[…]
@@
The algorithm $\mathsf{overlap}:\operatorname{\mathsf{Desc}}_{n}\times\operatorname{\mathsf{Desc}}_{n}\rightarrow\mathbb{C}$ runs in time $O(n^{3})$. It satisfies
$\mathsf{overlap}(\operatorname{\Delta}_{1},\operatorname{\Delta}_{2})$ $=\langle\psi(\operatorname{\Delta}_{1}),\psi(\operatorname{\Delta}_{2})\rangle$ (91)
for all $\operatorname{\Delta}_{1},\operatorname{\Delta}_{2}\in\operatorname{\mathsf{Desc}}_{n}\$.